# TREACLE: Thrifty Reasoning via Context-Aware LLM and Prompt Selection

### v1

@@ 5.2.2 Time-Varying API Query Latency
(a) Hard
(b) Easy
Figure 10: Performance of various methods on “easy” and “hard” partitions of the test set. The models are trained using the original training data, leading to a distribution shift in difficulty on the test set. Experiments with $\alpha=\frac{1}{20}$.
Table 3: Performance on “Easy” and “Hard” partitions of the test set. The total budget is \$$1$ and $\alpha=\frac{1}{20}$.
# of questions unanswered Budget spent (\$)
All Easy Hard All Easy Hard
TREACLE 1/1319 0/500 2/500 0.972 0.841 0.999
Single Model 2/1319 0/500 0/500 0.689 0.688 0.703
Cal. Cascade 2/1319 0/500 26/500 0.931 0.702 0.998
FrugalGPT 3/1319 0/500 31/500 0.955 0.713 0.998
5.2.3 Shifts in Question Difficulty
Thus far in the evaluations, easier and harder questions were evenly mixed throughout the training and test sets. In practice, easier and harder questions may not be uniformly spaced, and we would like TREACLE to be robust to that. In this subsection, we examine two types of non-uniform distributions of query difficulty: harder/easier problems in the training and test sets, and towards the end of the test set.
Difficulty shifts between training and test. We divided the GSM8K test set into “hard” and “easy” subsets based on the question difficulty. The difficulty is defined by the number of LLM models correctly answering the question. Intuitively, more models answering a question correctly means it is easier. Basic performance on the easy and hard questions is shown in Table 3. When the questions are hard, each question ends up consuming too much budget, leaving insufficient budget for subsequent questions that then go unanswered. The single model baseline does well in terms of cost and unanswered questions, but has low accuracy. We plot the performance for variable budgets in Figure 10, and find that TREACLE’s accuracy remains stable, no matter whether the test distribution shifts to an easier level or a harder level. This is because TREACLE can dynamically adjust based on the remaining budget in online fashion.
Figure 11: TREACLE is robust to re-ordered question difficulty in the test set. Experiments with $\alpha=\frac{1}{20}$.
5.2.4 Different types of reasoning tasks
We conducted experiments on various types of reasoning datasets to understand TREACLE’s performance, with full results given in the Appendix. To visualize the differences between the three datasets, in Figure 12 we plotted the fraction of questions where the most powerful (LLM, prompt) combination in the sorted list correctly answered the question (the “in order” pie slice), versus those questions where a less powerful combination succeeded and a more powerful combination failed (all other slices of the pie). Interestingly for all datasets, there are minority cases where less powerful LLMs (the smaller pieces of the pie) can answer the question correctly. Such cases are most prevalent in the GSM8K dataset and least prevalent in LLC, possibly because the math questions of GSM8K are more difficult. Despite these dataset differences, TREACLE still chooses the right (LLM, prompt) combination to achieve higher accuracy in all datasets than the baselines.
Refer to captionGSM8K
Refer to captionLLC
Figure 12: Fraction of questions that are solved by (LLM, prompt) combinations ordered from least to most powerful (“in order”). Minority slices are questions where less powerful combinations correctly answered.
Mixture of tasks. We also seek to examine whether one model can handle multiple types of tasks under one common budget. This is in contrast to the previous experiments where each task required a specialized model. Specifically, we trained a single model with all 3 datasets and recorded the test accuracy on the same 3 datasets. The results shown in Figure 13 for “TREACLE (all tasks)”, offline knapsack, and online knapsack are the test accuracy from an equal mix of CSQA, GSM8K and LLC queries. “TREACLE (individual tasks)” is the test accuracy on the same mix of queries, using the models from previous subsections, where each model (corresponding to a task) is assigned to 1/3 of the common budget. The results show that “TREACLE (all tasks)” can handle a mixture of tasks under a common budget (e.g., outperforming online knapsack), and can significantly outperform the individual tasks baseline (“TREACLE (individual tasks)”) by effectively allocating its common budget across queries of different types.
Figure 13: TREACLE can handle a mixture of tasks under a common budget.
Addition of an unseen new task type. We now consider the scenario where the model has not been trained on certain tasks. To show that TREACLE can adapt to new tasks easily, we performed additional experiments. The base model is trained using the CSQA dataset, and the unseen new tasks are queries from GSM8K. Interestingly, in our design, we decouple “decision making” from the “task embedding”: To transfer from CSQA to GSM8K, we freeze the base RL policy of CSQA (the decision making part), and fine-tune the “current question’s text embedding” feature in the state vector (the task embedding part) that feeds to this RL policy. The question is how many samples are needed to fine-tune the text embedding.
As shown in Figure 14, with a budget of 0.6, the original model fully-trained on GSM8K (“train on GSM8K”) achieves a test accuracy of 0.848, compared to 0.78 when trained on CSQA and fine-tuned with only 200 additional samples from GSM8K (“fine-tune on 200 GSM8K”). This highlights a relatively small accuracy loss when transferring to new types of unseen tasks. The results suggest that our method can seamlessly adapt to new tasks with only a small amount of additional training, which is important for overall cost savings.
Figure 14: Test accuracy on GSM8K
6 Conclusions
We proposed TREACLE, an learning-based LLM querying framework that intelligently chose between LLM and prompt combinations based on questoin context and past response history. Our experiments showed that TREACLE outperforms other baselines and is robust to different budgets, LLM availability and prices, query latency, etc. For future work, we plan to explore other non-reasoning tasks and incorporate other features such as privacy into the cost function. We hope our framework can help spur research on cost-efficient and robust LLM systems.
Broader impact. The potential broader impact of this work is to make LLMs cheaper to use and more accessible to cost-sensitive users.
Appendix A Implementation Details
During training, we used the Adam optimizer with a learning rate $41\times 10^{-}$ 4, Huber loss as the loss function, and a batch size of 64. Our DQN has three layers with ReLU and softmax activations, and the size of the hidden layer is 128. We set $\lambda=5$ in the reward function. For re-queries, we set different temperature settings for Llama and GPT (0.8 and 1, respectively) because their ranges are different ($[0,1]$ and $[0,2]$ respectively). The actions are selected according to an $\epsilon$-greedy policy. Simply put, the actions are sometimes chosen by the DQN and sometimes sampled uniformly. The probability of choosing a random action starts at $\epsilon_{\text{START}}=0.9$ and decays exponentially towards $\epsilon_{\text{END}}=0.05$.
Appendix B Understanding Optimal Cascade Strategy
Setup: Suppose there are $M$ models each with probability of correct answer $p_{i}$ and inference cost $c_{i}$ for $1\leq i\leq M$. During the cascade, we assume access to an oracle that tells when the answer of a model is incorrect so that we can move to the next model. The procedure continues until a correct answer is obtained. The goal is to minimize the expected cost of inference.
Lemma 1.
Suppose there are two models with probability of correct answers $p_{1}$ and $p_{2}$ and inference costs $c_{1}$ and $c_{2}$ respectively. Then, the optimal cascade rule is to first query the model with larger cost-normalized accuracy $p_{i}/c_{i}$.
Proof.
Cascade terminates when the first correct answer is obtained. The expected cost of inference if we query model 1 first is $c_{1}p_{1}+(c_{1}+c_{2})(1-p_{1})$. Note that accuracy is independent of the cascade order thanks to oracle. From this we wish to minimize $c_{i}p_{i}+(c_{1}+c_{2})(1-p_{i})=c_{1}+c_{2}-c_{1}c_{2}p_{i}/c_{i}$. This is minimized when $p_{i}/c_{i}$ is larger. ∎
Lemma 2.
Suppose there are $M$ models with probability of correct answer and cost pairs $(p_{i},c_{i})$ and suppose their outputs are independent of each other. Then, the optimal cascade rule is ordering models according to their cost-normalized accuracies $p_{i}/c_{i}$.
Proof.
Suppose the models are not ordered in terms of $p_{i}/c_{i}$. We will prove that swapping the order results in a better cascade. Again recall that accuracy is independent of the cost because we stop as soon as oracle confirms the answer. Expected accuracy is equal to the probability of at least one model being correct. To proceed, suppose models are not ordered and there is some $k$ such that $p_{k}/c_{k}<p_{k+1}/c_{k+1}$. Let us compute the expected change in inference cost when we flip their order.
The change in cost of inference arises from the scenarios model $k$ will be used but $k+1$ will not be and vice versa. Define $q_{k-1}=\prod_{i=1}^{k-1}(1-p_{i})$ which is the probability of first $k-1$ models failing. The probability of the excess cost associated to model $k$ then $k+1$ is
$ec_{k}=q_{k}c_{k}+q_{k}(1-p_{k})c_{k+1}.$
This is because if first $k-1$ models fail, we definitely pay $c_{k}$ and only pay $c_{k+1}$ if $k$ fails. This excess cost definition does not reflect the models strictly before $k$ or strictly after $k+1$. Because the expected cost of other models arise from the symmetric events with respect to models $k$ and $k+1$ (either problem is already solved by the time we reach $k$ or both $k,k+1$ failed).
Conversely, excess cost associated to $k+1$ is (i.e. using $k+1$ first and then $k$)
$ec_{k+1}=q_{k}c_{k+1}+q_{k}(1-p_{k+1})c_{k}.$
To proceed, observe that if $p_{k}/c_{k}<p_{k+1}/c_{k+1}$, then we would be better off by switching the models because
$ec_{k+1}<ec_{k}\iff$ (1)
$q_{k}c_{k+1}+q_{k}(1-p_{k+1})c_{k}<q_{k}c_{k}+q_{k}(1-p_{k})c_{k+1}\iff$ (2)
$c_{k+1}+(1-p_{k+1})c_{k}<c_{k}+(1-p_{k})c_{k+1}\iff$ (3)
$p_{k}c_{k+1}<p_{k+1}c_{k}\iff p_{k}/c_{k}<p_{k+1}/c_{k+1}.$ (4)
We also verify that our models are sorting following decreasing cost-normalized accuracies. The results for purely monetary pricing with $\alpha=\frac{1}{10}$ are shown in Figure 15. With other $\alpha$ values, Llama-2-13b will be cheaper, so the cost-normalized accuracy will always be decreasing.
Figure 15: Cost-normalized accuracy
Appendix C Additional Results
C.1 Model and Prompt Characterization
@@
C.2 Different cost function parameters and datasets
Providers may adjust the per-token API prices, or the user may value latency and monetary price differently. Therefore, we conducted experiments using different settings of the $\alpha$ (defined in Section 5.1.2)and $\beta$ (defined in Section 3) parameters in the cost function. In Figure 16, the cost ratio $\alpha$ increases from left to right, and hence the cost difference between more powerful (GPT) and weaker (Llama) models gradually decreases according to the definition. Under different pricing policies, ${\sf\small TREACLE}$ consistently achieves better performance than the online baselines. In other words, a single ${\sf\small TREACLE}$ model can easily accommodate varying budget requirements and cost functions, since it was trained under heterogeneous parameter settings.
@@ C.2 Different cost function parameters and datasets
Llama-2-7b “Follow example” simple CoT few-shot 23.36 23.65 909.81/911.43 120.49/119.14
@@
“Follow example” simple CoT few-shot 37.90 37.91 909.81/911.43 128.41/128.29
“Follow example” complex CoT few-shot 42.77 44.05 2943.81/2945.43 328.99/326.11
@@
“Follow example” simple CoT few-shot 82.00 79.15 772.05/773.70 107.23/108.31
“Follow example” complex CoT few-shot 83.30 82.94 2419.00/2416.00 82.00/49.00
@@
“Follow example” simple CoT few-shot 92.61 92.34 770.05/771.70 146.83/149.67
@@
“Follow example” simple CoT few-shot 93.95 92.95 770.05/771.70 101.51/103.62
@@
PaLM “Follow example” simple CoT few-shot 63.05 62.17 860.04/861.70 115.15/115.51
“Follow example” complex CoT few-shot 66.95 64.14 1918.04/1919.70 110.68/110.69
@@
“Follow example” standard few-shot 58.86 63.06 446.36/446.06 512.00/512.00
“Follow example” CoT few-shot 64.72 67.65 640.36/640.06 512.00/512.00
@@
“Follow example” standard few-shot 65.29 66.83 446.36/446.06 512.00/511.80
“Follow example” CoT few-shot 68.19 71.17 640.30/640.06 512.00/512.00
@@
“Follow example” standard few-shot 74.09 76.82 396.74/396.44 3.50/3.55
“Follow example” CoT few-shot 68.31 68.55 575.74/575.44 16.51/16.74
@@
“Follow example” standard few-shot 84.29 87.14 396.74/396.44 2.00/2.00
“Follow example” CoT few-shot 82.12 85.83 575.74/575.44 5.05/5.20
@@
“Follow example” standard few-shot 0.94 1.4 155.25/358.24 155.17/352.73
“Follow example” CoT few-shot 44.23 44.60 344.25/344.17 71.77/71.11
@@
“Follow example” standard few-shot 2.41 2.93 155.25/155.17 239.28/243.78
“Follow example” CoT few-shot 48.63 48.87 344.25/491.84 71.77/493.51
@@
“Follow example” standard few-shot 8.16 9.47 138.51/138.45 5.54/5.51
“Follow example” CoT few-shot 87.13 86.53 304.51/304.45 63.01/62.86
@@
“Follow example” standard few-shot 23.74 24.27 138.51/138.45 5.72/5.72
“Follow example” CoT few-shot 92.68 93.2 304.51/304.45 63.00/62.86
C.3 Ablation experiments
@@ Appendix E Full Prompts
Table 8: Domain expert prompting strategy (“Math solver” and “Math assistant”) in GSM8K dataset, where {Question} means that original question text.
@@
Question: Take the last letters of the words in “Elon Musk” and concatenate them.
@@
Question: Take the last letters of the words in “Larry Page” and concatenate them.
@@
Question: Take the last letters of the words in “Sergey Brin” and concatenate them.
@@
Question: Take the last letters of the words in ”Bill Gates” and concatenate them.
@@
Question: Take the last letters of the words in “Elon Musk” and concatenate them.
@@
The last letter of “Elon” is “n”.
The last letter of “Musk” is “k”.
Concatenating them is “nk”.
@@
Question: Take the last letters of the words in “Larry Page” and concatenate them.
@@
The last letter of ”Larry” is “y”.
The last letter of ”Page” is “e”.
Concatenating them is “ye”.
@@
Question: Take the last letters of the words in “Sergey Brin” and concatenate them.
@@
The last letter of “Sergey” is “y”.
The last letter of ”Brin” is “n”.
Concatenating them is “yn”.
@@
Question: Take the last letters of the words in ”Bill Gates” and concatenate them.
@@
The last letter of “Bill” is “l”.
The last letter of “Gates” is “s”.
Concatenating them is “ls”.

### v2

@@ C.1 Time-Varying API Query Latency
C.2 Model and Prompt Characterization
@@
C.3 Different types of reasoning tasks
To visualize the differences between the three reasoning datasets, in Figure 14 we plotted the fraction of questions where the most powerful (LLM, prompt) combination in the sorted list correctly answered the question (the “in order” pie slice), versus those questions where a less powerful combination succeeded and a more powerful combination failed (all other slices of the pie). Interestingly for all datasets, there are minority cases where less powerful LLMs (the smaller pieces of the pie) can answer the question correctly. Such cases are most prevalent in the GSM8K dataset and least prevalent in LLC, possibly because the math questions of GSM8K are more difficult. Despite these dataset differences, TREACLE still chooses the right (LLM, prompt) combination to achieve higher accuracy in all datasets than the baselines.
Refer to captionGSM8K
Refer to captionLLC
Figure 14: Fraction of questions that are solved by (LLM, prompt) combinations ordered from least to most powerful (“in order”). Minority slices are queries where less powerful combinations correctly answered.
C.4 Different cost function parameters and datasets
Providers may adjust the per-token API prices, or the user may value latency and monetary price differently. Therefore, we conducted experiments using different settings of the $\alpha$ (defined in Section A.2)and $\beta$ (defined in Section 3) parameters in the cost function. In Figure 15, the cost ratio $\alpha$ increases from left to right, and hence the cost difference between more powerful (GPT) and weaker (Llama) models gradually decreases according to the definition. Under different pricing policies, ${\sf\small TREACLE}$ consistently achieves better performance than the online baselines. In other words, a single ${\sf\small TREACLE}$ model can easily accommodate varying budget requirements and cost functions, since it was trained under heterogeneous parameter settings.
@@ C.4 Different cost function parameters and datasets
Refer to captionGSM8K, $\alpha=\frac{1}{50}$
Refer to captionGSM8K, $\alpha=\frac{1}{20}$
Refer to captionGSM8K, $\alpha=\frac{1}{10}$
Refer to captionGSM8K, $\beta$=1M
Refer to captionGSM8K, $\beta=$500k
Refer to captionGSM8K, $\beta=$50k
@@
Llama-2-7b “Follow example" simple CoT few-shot 23.36 23.65 909.81/911.43 120.49/119.14
@@
“Follow example" simple CoT few-shot 37.90 37.91 909.81/911.43 128.41/128.29
“Follow example" complex CoT few-shot 42.77 44.05 2943.81/2945.43 328.99/326.11
@@
“Follow example" simple CoT few-shot 82.00 79.15 772.05/773.70 107.23/108.31
“Follow example" complex CoT few-shot 83.30 82.94 2419.00/2416.00 82.00/49.00
@@
“Follow example" simple CoT few-shot 92.61 92.34 770.05/771.70 146.83/149.67
@@
“Follow example" simple CoT few-shot 93.95 92.95 770.05/771.70 101.51/103.62
@@
PaLM “Follow example" simple CoT few-shot 63.05 62.17 860.04/861.70 115.15/115.51
“Follow example" complex CoT few-shot 66.95 64.14 1918.04/1919.70 110.68/110.69
@@
“Follow example" standard few-shot 58.86 63.06 446.36/446.06 512.00/512.00
“Follow example" CoT few-shot 64.72 67.65 640.36/640.06 512.00/512.00
@@
“Follow example" standard few-shot 65.29 66.83 446.36/446.06 512.00/511.80
“Follow example" CoT few-shot 68.19 71.17 640.30/640.06 512.00/512.00
@@
“Follow example" standard few-shot 74.09 76.82 396.74/396.44 3.50/3.55
“Follow example" CoT few-shot 68.31 68.55 575.74/575.44 16.51/16.74
@@
“Follow example" standard few-shot 84.29 87.14 396.74/396.44 2.00/2.00
“Follow example" CoT few-shot 82.12 85.83 575.74/575.44 5.05/5.20
@@
“Follow example" standard few-shot 0.94 1.4 155.25/358.24 155.17/352.73
“Follow example" CoT few-shot 44.23 44.60 344.25/344.17 71.77/71.11
@@
“Follow example" standard few-shot 2.41 2.93 155.25/155.17 239.28/243.78
“Follow example" CoT few-shot 48.63 48.87 344.25/491.84 71.77/493.51
@@
“Follow example" standard few-shot 8.16 9.47 138.51/138.45 5.54/5.51
“Follow example" CoT few-shot 87.13 86.53 304.51/304.45 63.01/62.86
@@
“Follow example" standard few-shot 23.74 24.27 138.51/138.45 5.72/5.72
“Follow example" CoT few-shot 92.68 93.2 304.51/304.45 63.00/62.86
C.5 Ablation experiments
@@ Appendix E Full Prompts
Table 7: Domain expert prompting strategy (“Math solver" and “Math assistant") in GSM8K dataset, where {Question} means that original question text.
@@
Question: Take the last letters of the words in “Elon Musk" and concatenate them.
@@
Question: Take the last letters of the words in “Larry Page" and concatenate them.
@@
Question: Take the last letters of the words in “Sergey Brin" and concatenate them.
@@
Question: Take the last letters of the words in "Bill Gates" and concatenate them.
@@
Question: Take the last letters of the words in “Elon Musk" and concatenate them.
@@
The last letter of “Elon" is “n".
The last letter of “Musk" is “k".
Concatenating them is “nk".
@@
Question: Take the last letters of the words in “Larry Page" and concatenate them.
@@
The last letter of "Larry" is “y".
The last letter of "Page" is “e".
Concatenating them is “ye".
@@
Question: Take the last letters of the words in “Sergey Brin" and concatenate them.
@@
The last letter of “Sergey" is “y".
The last letter of "Brin" is “n".
Concatenating them is “yn".
@@
Question: Take the last letters of the words in "Bill Gates" and concatenate them.
@@
The last letter of “Bill" is “l".
The last letter of “Gates" is “s".
Concatenating them is “ls".